Hot combustion gases of a furnace are flowing through a square chimney made of concrete (k = 1.4 W/m-K). The flow section of the chimney is 20 cm x 20 cm, and the thickness of the wall is 20 cm. The average temperature of the hot gases in the chimney is T, = 300°C, and the average convection heat transfer coefficient inside the chimney is h; = 70 W/m-K. The chimney is los ing heat from its outer surface to the ambient air at T = 20°C by convection with a heat transfer coefficient of h,= 21 W/m2K and to the sky by radiation. The emissivity of the outer surface of the wall is e=0.9, and the effective sky temperature is estimated to be 260 K. Using the finite difference method with Ax = Ay = 10 cm and taking full advantage of symmetry, determine the tem pressure at the nodal points of a cross section and the rate of heat loss for a 1-m-long section of the chimney.
1
Expert's answer
2020-09-22T17:04:43-0400

Given data- this question basically related to conduction, convection, Rediation.

Hot side temp.T1 =300+273= 573K

Thermal conductive heat transfer coefficient K=1.4 w/m-k

Convective heat transfer coefficient h1= 70w/ m2-k

Convective heat transfer coefficient h2 =21w/m2-k

Area of chimney "= 20\times 20" cm2

Thickness t = 20 cm

Emissivity of outer surface e = 0.9

Sky temp. = 260 k

1- Temp. At nodal point = ?

2- "\frac{dQ}{dL} = ?" = ?

"dQ" "= KA\frac{(T1-T2)}{L}"